OpenStudy (anonymous):

Write a quadratic equation having the given numbers as solutions. -10 and -1 The quadratic equation is _ = 0

OpenStudy (anonymous):

(x+10)(x+1) then factor

OpenStudy (anonymous):

(x+10)(x+1)=0 x^2+10x+x+10=0 x^2+11x+10=0